OpenStudy (anonymous):

What is the volume of the shaded region? Round your answer to 2 decimal places if necessary. Use π = 3.14 when necessary

OpenStudy (anonymous):

whats the shaded region????

OpenStudy (anonymous):

volume of prism - volume of sphere

OpenStudy (anonymous):

The volume of the cube is 103=1000 , the sphere has radius 5, so it's volume is 4/3∗π∗r3=4/3∗3.14∗53=523.33 The volume of the shaded region is the volume of the cube minus the volume of the sphere so 1000−523.33=476.67

OpenStudy (anonymous):

thanx